A large, flat, horizontal sheet of charge has a charge per unit area of 4.70 μC/m2. Find the electric field just above the middle of the sheet. magnitude direction upward kN/C
Answer #1

E = σ/(2єₒ ) σ =4.70 * (10^ -6) , єₒ=8.854*(10^-12)

E= (4.70*10^ -6)/(2 x 8.85 X 10^-12) =0.265* 10^6= 265 kN/C